2. Two chambers A and B are separated by a
membrane which is permeable to Cl⁻ , but
not to K⁺. Chamber A is filled with 100 mM
KCI and chamber B with 10mM KCl. For
the above situation answer the following
questions :

(a) Do these chambers remain in state of
electrical neutrality before any
movement of the ions starts ? Give
reasons.

(b) Explain what happens when the
diffusion starts.
1
Expert's answer
2019-11-14T09:33:34-0500

(a)Yes,these chambers remain in a state of electrical neutrality as there are same number of positive ions as there are negative ions.

(b)When diffusion starts,as concentration of KCl in chamber A is more than the concentration of KCl in chamber B,so,Cl-moves from Chamber A to Chamber B.This diffusion takes place until they attain equilibrium i.e concentration in Chamber A is equal to concentration in Chamber B.

10×V+100×V=x×2V

10×V+100×V=x×2V

x=110/2=55mM

x=110/2=55mM (Answer)

This is the equilibrium concentration.